# Choosing observables that capture critical slowing down before tipping points: A Fokker-Planck operator approach

### v1

@@ I Introduction
We consider how critical transitions in stochastically forced complex systems may be anticipated by measuring increases in amplitude and temporal correlation of fluctuations in certain observables as early-warning signals (EWS). We consider a general class of heterogeneous systems with many interacting agents or scales, arising for instance in ecology, biology, social science, and the Earth system [1]. Such complex systems are often modeled by a first-order stochastic differential equation with a non-linear, deterministic drift giving rise to possibly chaotic dynamics, and a noise process that represents unresolved scales and random disturbances by the environment, as well as control parameters that represent slow changes in external boundary conditions. A critical transition occurs when upon change of a control parameter a base state, i.e., a stable invariant set of the drift, loses stability and the system undergoes an abrupt transition to an alternative state. This is usually due to a collision of the base state with an edge state, which is an unstable invariant set of the drift. The stable manifold of the edge state is the basin boundary separating the base state from the alternative state. The simplest case of such a transition is a noisy saddle-node bifurcation (SNB) [2], which is often considered the archetype of a tipping point (TP).
@@
EWS arise due to so-called critical slowing down (CSD) [3, 4, 5, 6, 7, 8, 9], which is most easily understood for the SNB. Here, as a control parameter $\mu$ that modulates the drift crosses a critical value (e.g. $\mu_{c}=0$) the leading eigenvalue of the Jacobian describing the linearized dynamics around the base fixed point crosses the imaginary axis. Leading up to this, for $0<\mu\ll 1$, the dynamics along one degree of freedom (d.o.f) becomes much slower compared to all others, and after a short relaxation time the system is confined to an (extended) center manifold, or more precisely to a neighborhood thereof due to noise from random perturbations of the system’s environment [10]. The drift on the manifold is one-dimensional and given by the SNB normal form after a suitable coordinate transformation. As $\mu\to 0$, CSD refers to the slowing down of the relaxation dynamics towards the equilibrium along the center manifold after an arbitrary, discrete perturbation.
@@
In more general cases, where the base state is a limit cycle or chaotic attractor, we also expect a decrease of the relaxation rate back to steady state after a perturbation of the system. This is plausible if we assume that upon control parameter change the underlying deterministic dynamics experience a continuous change from being stable to neutrally stable in one d.o.f, before finally becoming unstable. This generic feature makes the detection of loss of resilience to perturbations the primary avenue for predicting TPs [11].
@@
For real-world, large-scale systems controlled perturbations are not available. Instead, there is a permanent influence of random disturbances from the environment. Such noise-driven, natural fluctuations of the unperturbed system allow one to infer the system’s response to perturbations if linear response theory guarantees a fluctuation-dissipation theorem [12, 13]. In particular, the size and correlation of the fluctuations are expected to grow in tandem with the system’s slowing response as the critical transition is approached - which is the other side of the coin of CSD - thereby forming statistical EWS. Growing fluctuations towards the basin boundary imply a flattening of the quasipotential [14]. This happens in the direction of a particular d.o.f that is related to the location of the edge state, since the latter usually lies on the most probable path of a noise-induced escape [15].
@@
Real-world observations have been analyzed for CSD by measuring statistical EWS of presumed critical transitions, including financial crises [16], depression [17, 18], neuron spiking [19], and climate tipping points, such as the Greenland ice sheet [20], Amazon forest [21] and Atlantic Meridional Overturning Circulation (AMOC) [22, 23, 24]. But statistical false positives and false negatives can occur. The destabilization of the system in a single (critical) d.o.f implies that the increase in noise-driven variability occurs also predominantly in a single d.o.f, which gives rise to a scalar observable where the variance is expected to diverge and the autocorrelation to tend to 1 at the bifurcation. Thus, EWS can be masked if the measurements have been taken from a dynamical observable that does not sufficiently project on the critical d.o.f [2, 25, 26, 27, 15]. In the simple case of the SNB, this means that the observable does not follow the SNB normal form to any good approximation and hence does is not subjected to significant CSD.
@@
Consequently, the central question that will be addressed here is what observables should be used to detect CSD. This depends on how the system under the influence of noise responds to perturbations away from its steady state, and how this response changes as a control parameter slowly approaches its critical value. This can be understood in terms of the Fokker-Planck (FP) equation, which governs the temporal evolution of the probability density of the state in phase space. The density can be written as an expansion in the eigenfunctions $\psi_{n}$ of the FP operator $\mathcal{L}$. For a fixed control parameter any initial density will converge to the unique stationary density $\pi(\mathbf{x})$, which is the first eigenfunction $\psi_{0}$ with eigenvalue $\lambda_{0}=0$. The system is then in statistical equilibrium, where the contributions of all other $\psi_{n}$ with $\lambda_{n}<0$ have decayed. The first few $\psi_{n}$ (with $\lambda_{n}$ closest to 0) signify locations in phase space where fluctuations tend to linger on a finite, but long-term time horizon.
@@
We consider systems with a TP, i.e., the deterministic drift is (at least) bistable with a base and alternative attractor. In accordance with the paradigm of bifurcation-induced tipping [28] we assume low noise, and hence the system spends long periods of time in distinct regions around the attractors of the drift, referred to as metastable states. There are rare transitions between the states on time scales of $\mathcal{O}(1/\lambda_{1})$, and the eigenfunction $\psi_{1}$ with $|\lambda_{1}|\ll 1$ signifies a very slow transfer of density between the metastable states. While part of the invariant density $\pi(\mathbf{x})$ occupies the alternative metastable state, we can assume that on a finite time horizon the system is in a quasi-stationary distribution concentrated entirely around the base state, where the contribution of $\psi_{1}$ has not decayed. Due to CSD, the relaxation towards the base state within the quasipotential well of the base state becomes slower along a particular mode. When close enough to the TP, this mode becomes the slowest in the system and will be expressed by the next eigenfunction $\psi_{2}$.
@@
An observable that naturally expresses increases in fluctuations related to this critical mode is proposed here to be given by the corresponding eigenfunction $\phi_{2}$ of the backward (adjoint) operator $\mathcal{L}^{*}$, also known as the generator, which shares the same eigenvalues as $\mathcal{L}$. The latter governs the temporal evolution of expectation values of observables as a function of initial states, and the first few $\phi_{n}$ can be interpreted as patterns of initial conditions with slowest decay towards $\pi(\mathbf{x})$. This proposition is in agreement with the framework of optimal fingerprints presented in [29].
@@
To obtain $\mathcal{L}^{*}$ from observational data we propose to use the diffusion map (DM) algorithm [30, 31, 32]. DM has been successfully used to define generalized collective coordinates that capture the effective dynamics of complex systems [32, 33, 34]. It gives an approximation (discretized on the set of data points) of $\mathcal{L}^{*}$ induced by a stochastic differential equation with drift $\nabla\ln[\pi(x)]$, i.e., a gradient system related to the quasipotential of the underlying stochastic dynamic system [14, 35]. This incurs an error when the underlying system has strong non-gradient dynamics, but it should still give useful results in our context since it preserves the flattening of the quasipotential (in the direction of a particular critical d.o.f) as a key property of CSD, which is not affected by non-gradient terms of the drift. Indeed, we show that from a DM approximation we can obtain physical observables that carry excellent EWS also for non-gradient systems, including a high-dimensional global ocean model exhibiting a TP of the AMOC.
@@
The paper is structured as follows. Sec. II reviews some fundamentals of FP operators and its spectral properties and introduces notation, as well the DM algorithm to estimate the eigenfunctions of $\mathcal{L}^{*}$ from data. In Sec. III.1 and III.2 we motivate the usage of the backward FP eigenfunctions for the purposes of EWS with simple double well potential systems in one and two dimensions. In Sec. III.3 we show that the reconstruction of the eigenfunctions with the DM method indeed yields observables that carry strong EWS in conceptual models. Further, in Sec. III.4 we show that such observables are strictly necessary if one wants extrapolate from increasing fluctuations to forecast the timing of a tipping point. In Sec. III.5 we apply our method successfully to a high-dimensional model of the global ocean circulation, and conclude with a discussion in Sec. IV.
@@ II Fokker-Planck eigenfunctions and diffusion maps
Consider the $d$-dimensional system state $\mathbf{X}=(X^{1},...,X^{d})$, governed by an Ito diffusion equation with time-independent coefficients written component-wise as
$dX^{\gamma}=a^{\gamma}(\mathbf{X})dt+\sqrt{\epsilon}\sigma_{\nu}^{\gamma}(\mathbf{X})dW^{\nu},$ (1)
with drift $a^{\gamma}(\mathbf{X})$ and diffusion $\sigma_{\nu}^{\gamma}(\mathbf{X})$. The transition probability density $P(\mathbf{X}(t)=\mathbf{x}|\mathbf{X}(0)=\mathbf{x_{0}})\equiv P(\mathbf{x},t|\mathbf{x_{0}})$ is governed by the Fokker-Planck (FP) equation
@@
$\mathcal{L}(x)=-\frac{\partial}{\partial x^{\gamma}}a^{\gamma}(\mathbf{x})+\frac{\epsilon}{2}\frac{\partial^{2}}{\partial x^{\gamma}\partial x^{\nu}}b^{\gamma\nu}(\mathbf{x})$ (3)
and diffusion tensor $b^{\gamma\nu}(\mathbf{x})=\sigma_{\lambda}^{\gamma}(\mathbf{x})\sigma_{\sigma}^{\nu}(\mathbf{x})\delta^{\lambda\sigma}$. The stationary distribution $\pi(\mathbf{x})\equiv\lim_{t\to\infty}P(\mathbf{x},t|\mathbf{x_{0}})$ satisfies $\mathcal{L}\pi=0$ and is an eigenfunction of $\mathcal{L}$ with eigenvalue $\lambda_{0}=0$.
We assume in the following that the system obeys detailed balance, which is the case in gradient systems with additive noise. As mentioned above, we still apply our method to non-gradient systems, observing that we will effectively reconstruct a gradient system based on the quasipotential $V_{q}(\mathbf{x})\propto\ln\pi(\mathbf{x})$ of the full system, and that the presence of non-gradient terms does not change $\pi(\mathbf{x})$. The eigenfunctions of the FP operator
@@
with eigenvalues $\lambda_{0}=0>\lambda_{1}\geq\lambda_{2}\geq...>-\infty$ then form an orthonormal basis under an inner product weighted by $\pi(\mathbf{x})$. Time-varying solutions of the FP equation can be written in the eigenfunction basis as
@@
[…]
@@
$-\partial_{s}u(\mathbf{x},s)=\mathcal{L}^{*}u(\mathbf{x},s),$ (7)
which is defined on the time interval $s\in[0,T]$ for functions $u(\mathbf{x},s)=\mathbb{E}_{x,s}[f(X_{T})]\equiv\mathbb{E}[f(X_{T})|X_{s}=\mathbf{x}]$, i.e., conditional expectation values of observables $f(\mathbf{x},s)$ (initialized at $\mathbf{x}$), and with the final condition $u(x,T)=f(x)$. Employing the transformation $t=T-s$ the equation can be formulated as an initial value problem
@@
[…]
@@
For a (large) data set, a discrete approximation to the (first few) $\phi_{n}(x)$ can be obtained by the diffusion map (DM) algorithm. We briefly summarize the method, and for more details refer the reader to, e.g., [36]. The algorithm defines a weighted graph on the data points, and subsequently computes the first few eigenvalues and eigenvectors of a random walk on this graph. To this end we define a kernel with bandwidth $\epsilon>0$ measuring the distance of two data points $\mathbf{x}$ and $\mathbf{y}$
$K(\mathbf{x},\mathbf{y})=\exp(-||\mathbf{x}-\mathbf{y}||^{2}/\epsilon^{2}).$ (11)
@@
with $D_{i}=\sum_{j=1}^{N}\tilde{K}_{ij}$. The first few eigenvectors $\nu_{n}$ of $M$, corresponding to eigenvalues $\lambda_{n}^{(M)}$, define the so-called diffusion coordinates $\xi_{n}=\lambda_{n}^{(M)}\nu_{n}$. In the limit $N\to\infty$ and $\epsilon\to 0$, the operator $(M-I)/\epsilon$ converges to the adjoint $\mathcal{L}^{*}$ of the FP operator (i.e. the backward FP operator), and $\lambda_{n}^{(M)}\to\lambda_{n}$ [31, 37, 38]. The Euclidean distance between data points in the DM coordinates is called the diffusion distance. The diffusion distance measures how closely two points are connected via diffusion of the Markov chain $M$. Two points $\mathbf{x}$ and $\mathbf{y}$ may have small Euclidian distance, but large diffusion distance, which can reflect that the dynamics evolve on a lower dimensional manifold.
To evaluate the eigenfunctions approximately at points $\mathbf{y}$ that are not part of the given data set $\{\mathbf{x}_{i}\}_{i=1}^{N}$, we employ the so-called Nyström eigenspace interpolation
@@
and $D(\mathbf{y})=\sum_{i=1}^{N}\tilde{K}(\mathbf{y},\mathbf{x}_{i})$. In Eq. 14, $\xi_{n}(\mathbf{x}_{i})$ denotes the entry of the n-th eigenvector corresponding to the i-th data point.
In our implementation of the DM algorithm we normalize each data variable to have unit variance. Furthermore, we remove a small number of single and double outliers so that we can use a smaller $\epsilon$ to obtain a better approximation of the backward eigenfunctions. Single outliers are those data points that have the largest distance to its nearest data point, i.e., where the minimum of the distance to all other points is largest. We find $n=15$ points with the largest minimum distance and remove the respective columns and rows in the distance matrix $K(\mathbf{x}_{i},\mathbf{x}_{j})$ in (12). Thereafter we remove $n=10$ double outliers, which are those pairs of points where the second smallest distance to all other points is largest. The number of removed outliers has been chosen by trial and error to give the best performance across all data sets used here, which have a typical sample size of about 10,000. If this step is skipped, often a quite large $\epsilon$ is needed to prevent the first diffusion coordinates from not merely acting to cluster individual outliers against the rest of the data.

### v2

@@ I Introduction
We consider how critical transitions in stochastically forced complex systems may be anticipated by measuring increases in amplitude and temporal correlation of fluctuations in certain observables as early-warning signals (EWS). The complex systems of interest are a general class of heterogeneous systems with many interacting agents or scales, arising for instance in ecology, biology, social science, and the Earth system [1]. They are often modeled by a first-order stochastic differential equation with a non-linear, deterministic drift giving rise to possibly chaotic dynamics, and a noise process that represents unresolved scales and random disturbances by the environment, as well as control parameters that modify the drift and represent slow changes in external boundary conditions. A critical transition occurs when upon a parameter change a base state, i.e., a stable invariant set of the drift, loses stability and the system undergoes an abrupt transition to an alternative state. This is usually due to a collision of the base state with an edge state, which is an unstable invariant set of the drift. The stable manifold of the edge state is the basin boundary separating the base state from the alternative state. The simplest case of such a transition is a noisy saddle-node bifurcation (SNB) [2], which is often considered the archetype of a tipping point (TP).
@@
EWS arise due to so-called critical slowing down (CSD) [3, 4, 5, 6, 7, 8, 9], which is most easily understood for the SNB. Here, as a control parameter $\mu$ crosses a critical value (e.g. $\mu_{c}=0$), the leading eigenvalue of the Jacobian describing the linearized dynamics around the base fixed point crosses the imaginary axis. Leading up to this, for $0<\mu\ll 1$, the dynamics along one degree of freedom (d.o.f) becomes much slower compared to all others, and after a short relaxation time the system is confined to an (extended) center manifold, or a neighborhood thereof due to the noise forcing [10]. The drift on the manifold is one-dimensional and given by the SNB normal form after a suitable coordinate transformation. As $\mu\to 0$, CSD refers to the slowing down of the relaxation dynamics towards the equilibrium along the center manifold after an arbitrary perturbation.
@@
In more general cases, where the base state is a limit cycle or chaotic attractor, we also expect a decrease of the relaxation rate back to steady state after a perturbation of the system. This is plausible since upon control parameter change the underlying deterministic dynamics experience a continuous change from being stable to neutrally stable in one d.o.f, before finally becoming unstable. This generic feature makes the detection of loss of resilience to perturbations the primary avenue for predicting TPs [11].
@@
For large-scale systems in the real world, controlled perturbations are difficult to find. But there is a permanent influence of random disturbances from the environment. Such noise-driven, natural fluctuations of the unperturbed system allow one to infer the system’s response to perturbations if linear response theory guarantees a fluctuation-dissipation theorem [12, 13]. The size and correlation of the fluctuations are expected to grow in tandem with the system’s slowing response as the critical transition is approached, thereby forming statistical EWS. This is the other side of the coin of CSD. Growing fluctuations towards the basin boundary imply a flattening of the quasipotential [14]. This happens in the direction of a particular d.o.f that is related to the location of the edge state, since the latter usually lies on the most probable path of a noise-induced escape [15].
@@
Real-world observations have been analyzed for CSD by measuring statistical EWS of presumed critical transitions, including financial crises [16], depression [17, 18], neuron spiking [19], and climate tipping points, such as the Greenland ice sheet [20], Amazon forest [21] and Atlantic Meridional Overturning Circulation (AMOC) [22, 23, 24]. But statistical false positives and false negatives can occur. The destabilization of the system in a single (critical) d.o.f implies that the increase in noise-driven variability occurs also predominantly in a single d.o.f, which gives rise to a scalar observable where the variance is expected to diverge and the autocorrelation to approach 1 at the bifurcation. Thus, EWS can be masked if measurements have been taken from a dynamical observable that does not sufficiently project on the critical d.o.f [2, 25, 26, 27, 15]. In the simple case of the SNB, this means that the observable does not follow the SNB normal form to any good approximation and hence is not subjected to significant CSD.
@@
Consequently, the central question that will be addressed here is what observables should be used to detect CSD. This depends on how the system under the influence of noise responds to perturbations away from its steady state, and how this response changes as a control parameter approaches its critical value. This can be understood in terms of the Fokker-Planck (FP) equation, which governs the temporal evolution of the probability density of the state in phase space. The density can be written as an expansion in the eigenfunctions $\psi_{n}$ of the FP operator $\mathcal{L}$. For a fixed control parameter any initial density will converge to the unique stationary density $\pi(\mathbf{x})$, which is the first eigenfunction $\psi_{0}$ with eigenvalue $\lambda_{0}=0$. The system is then in statistical equilibrium, where the contributions of all other $\psi_{n}$ with $|\lambda_{n}|>0$ have decayed. The first few $\psi_{n}$ (with $|\lambda_{n}|$ closest to 0, and assuming a discrete spectrum) signify locations in phase space where fluctuations tend to linger on finite, but long-term time horizons.
@@
We consider systems with a possible TP, i.e., the deterministic drift term in the governing equations gives rise to (at least) bistability, with a base and an alternative attractor. A low noise strength is assumed, meaning it is small compared to the (quasi-)potential barrier height between competing attractors of the drift. It is assumed to hold for most values of $\mu$, except when very close to the bifurcation where the barrier height goes to zero. This is required for the paradigm of bifurcation-induced tipping [28] to be meaningful and for the concept of EWS to be useful. Hence, when far from the bifurcation, the system spends very long periods of time in distinct regions around the attractors, referred to as metastable states. Transitions between the states can be ignored, as they occur very rarely on time scales of $\mathcal{O}(1/|\lambda_{1}|)$ with $|\lambda_{1}|\ll 1$. While part of the invariant density $\pi(\mathbf{x})$ occupies the alternative metastable state, we hence assume that on a finite time horizon the system is in a quasi-stationary distribution concentrated entirely around the base state, where the contribution of $\psi_{1}$, which signifies a very slow transfer of density between the metastable states, has not decayed. Due to CSD, the relaxation towards the base state within the quasipotential well of the base state becomes slower along a particular mode. When close enough to the TP, this mode becomes the slowest in the system and will be expressed by the next eigenfunction $\psi_{2}$.
@@
An observable that naturally expresses increases in fluctuations related to this critical mode is proposed here to be given by the corresponding eigenfunction $\phi_{2}$ of the backward (adjoint) operator $\mathcal{L}^{*}$, also known as the generator. $\mathcal{L}^{*}$ governs the temporal evolution of expectation values of observables as a function of initial states (e.g. states after a perturbation), and the first few $\phi_{n}$ can be interpreted as patterns of initial conditions with slowest decay towards $\pi(\mathbf{x})$. This proposition is in agreement with the framework of optimal fingerprints presented in [29].
@@
To obtain $\mathcal{L}^{*}$ from data we propose to use the diffusion map (DM) algorithm [30, 31, 32]. DM has been successfully used to define generalized collective coordinates that capture the effective dynamics of complex systems [32, 33, 34]. It gives an approximation (discretized on the set of data points) of $\mathcal{L}^{*}$ induced by a stochastic differential equation with drift $\nabla\ln[\pi(x)]$, i.e., a gradient system related to the quasipotential of the underlying stochastic dynamic system [14, 35]. We show here that it yields physical observables carrying excellent EWS, also for non-gradient systems including a high-dimensional global ocean model exhibiting a TP of the AMOC. It preserves the flattening of the quasipotential (in a particular critical d.o.f) as a key property of CSD, which is not affected by non-gradient terms of the drift. Nevertheless, methods to approximate $\mathcal{L}^{*}$ for non-gradient systems can be applied, such as the extension of DM for general diffusions with multiplicative noise [36], which is also implemented here and shown to give very similar results (App. A).
@@
The paper is structured as follows. Sec. II reviews some fundamentals of FP operators and introduces notation, as well the DM algorithm to estimate the eigenfunctions of $\mathcal{L}^{*}$ from data. In Sec. III.1 and III.2 we motivate the usage of the backward FP eigenfunctions for the purposes of EWS with simple double well potential systems in one and two dimensions. In Sec. III.3 we show with conceptual models that the reconstruction of the eigenfunctions with DM indeed yields observables that carry strong EWS. Further, in Sec. III.4 we show that such observables are strictly necessary if one wants extrapolate from increasing fluctuations to forecast the timing of a TP. In Sec. III.5 we apply our method successfully to a high-dimensional model of the global ocean circulation, and conclude with a discussion in Sec. IV.
@@ II Fokker-Planck eigenfunctions and diffusion maps
Consider the $d$-dimensional state $\mathbf{X}=(X^{1},...,X^{d})$, governed by an Ito diffusion equation with time-independent coefficients written component-wise as
$dX^{\gamma}=b^{\gamma}(\mathbf{X})dt+\sqrt{\epsilon}\sigma_{\nu}^{\gamma}(\mathbf{X})dW^{\nu},$ (1)
with drift $b^{\gamma}(\mathbf{X})$ and diffusion $\sigma_{\nu}^{\gamma}(\mathbf{X})$. The transition probability density $P(\mathbf{X}(t)=\mathbf{x}|\mathbf{X}(0)=\mathbf{x_{0}})\equiv P(\mathbf{x},t|\mathbf{x_{0}})$ is governed by the Fokker-Planck (FP) equation
@@
$\mathcal{L}(x)=-\frac{\partial}{\partial x^{\gamma}}b^{\gamma}(\mathbf{x})+\frac{\epsilon}{2}\frac{\partial^{2}}{\partial x^{\gamma}\partial x^{\nu}}A^{\gamma\nu}(\mathbf{x})$ (3)
and diffusion tensor $A^{\gamma\nu}(\mathbf{x})=\sigma_{\lambda}^{\gamma}(\mathbf{x})\sigma_{\sigma}^{\nu}(\mathbf{x})\delta^{\lambda\sigma}$. The stationary distribution $\pi(\mathbf{x})\equiv\lim_{t\to\infty}P(\mathbf{x},t|\mathbf{x_{0}})$ is an eigenfunction of $\mathcal{L}$ with eigenvalue $\lambda_{0}=0$, satisfying $\mathcal{L}\pi=0$. The eigenfunctions of the FP operator
@@
have eigenvalues with $|\lambda_{0}|=0<|\lambda_{1}|\leq|\lambda_{2}|\leq$$\infty$ and negative real parts. We restrict our attention to a discrete FP operator spectrum. For certain sufficiently chaotic, deterministic dynamics $\mathcal{L}$ is quasi-compact in appropriate function spaces, which implies that the dominant eigenvalues are discrete and well separated from the essential spectrum [37]. Further, sufficiently smooth stochastic dynamical systems have compact operators provided the domain is compact [38], or in unbounded domains given certain conditions on the drift [39]. We only consider stochastic dynamical systems, since otherwise the concept of EWS is not meaningful. This can also be motivated physically in terms of the Hasselmann program [40], where stochastic dynamical systems are viewed as the limiting equations for macroscopic slow processes in high-dimensional deterministic systems. In particular, the interaction between chaotic multi-scale microscopic constituents in deterministic high-dimensional systems leads to an emergent stochastic behaviour at the slow macroscopic level (see, for example, [41]). A bounded domain may be physically justified too, since for practical purposes in reasonable time horizons the dynamics should be sufficiently contained to access only a bounded region.
Time-varying solutions of the FP equation can be written in the eigenfunction basis as
@@
with $c_{n}=\int\psi_{n}(\mathbf{x})\pi^{-1}(\mathbf{x})\pi_{0}(\mathbf{x})dx$, where $\pi_{0}(\mathbf{x})=P(\mathbf{x},t=0)$. Since $|\lambda_{0}|=0<|\lambda_{1}|\leq|\lambda_{2}|\leq$... with nonpositive real parts, eigenfunctions with small indices decay slowest. In view of studying TPs, we consider here dynamical systems that spend a very long time in one part of phase space (a metastable set) before exhibiting a rare transition to another, and so on. In the simplest case there are two metastable sets, corresponding to neighborhoods of the attractors of the underlying deterministic dynamics, which is reflected in the spectrum of $\mathcal{L}$ as $|\lambda_{1}|\ll 1$ and a spectral gap with $|\lambda_{2}-\lambda_{1}|\gg|\lambda_{1}|$. To study the dynamics before the TP, it is sufficient to consider this bistable case, since a critical transition generically consists of the collision of only a single base state and a single boundary.
@@
$\partial_{s}u(\mathbf{x},s)=\mathcal{L}^{*}u(\mathbf{x},s),$ (7)
which is defined on the time interval $s\in[0,T]$ for functions $u(\mathbf{x},s)=\mathbb{E}_{x,s}[f(X_{T})]\equiv\mathbb{E}[f(X_{T})|X_{s}=\mathbf{x}]$, i.e., conditional expectation values of observables $f(\mathbf{x},s)$ (initialized at $\mathbf{x}$), and with the final condition $u(\mathbf{x},T)=f(\mathbf{x})$. Employing the transformation t=Tst=T-s the equation can be formulated as an initial value problem
@@
with $d_{n}=\int\phi_{n}(\mathbf{x})f(\mathbf{x})\pi(\mathbf{x})d\mathbf{x}$, and eigenfunctions $\phi_{n}$ satisfying $\mathcal{L}^{*}\phi_{n}=\lambda_{n}\phi_{n}$. The leading eigenfunction $\phi_{0}(\mathbf{x})$, corresponding to the eigenvalue $\lambda_{0}=0$, is the unique solution of $\mathcal{L}^{*}\phi_{0}=0$ and is given by $\phi_{0}=\text{const}$. This reflects the ergodicity of the underlying system, which implies that (long-term) expectation values do not depend on the initial conditions.
@@
and possibly approximated by a truncation thereof. Here, $g_{n}=\int g(x)\phi_{n}(x)\pi(x)dx$. Thus, $\phi_{n}(\mathbf{x})$ can themselves be considered as observables, and in particular the first few $\phi_{n}(\mathbf{x})$ are special observables with expectation values that converge only slowly because they are non-constant functions in regions of phase space where there is a slow relaxation to equilibrium. From a different point of view, the subset of leading $\phi_{n}(x)$ are transformations of the system from the original coordinates to reduction coordinates. The reduction is meaningful in case of time scale separation, which is expected to emerge when the deterministic drift of the system approaches a bifurcation. In particular, it can be shown that the evolution of the first $k$ eigenfunctions is approximately Markovian [42]. In this case, the long-term evolution of the system is governed by the first $k$ backward eigenfunctions.
For systems obeying detailed balance, i.e., when the drift $b(\mathbf{X})$ is the gradient of a potential and the diffusion $\sigma(\mathbf{X})$ is independent of the position $\mathbf{X}$ (additive noise), a discrete approximation to $\phi_{n}(x)$ can be obtained by the diffusion map (DM) algorithm given sufficient data. For the majority of this work, we will use this method and show that it gives good results in the context of EWS, even for non-gradient systems, where we thus effectively reconstruct a gradient system based on the quasipotential $V_{q}(\mathbf{x})\propto\ln\pi(\mathbf{x})$ of the full system. The presence of non-gradient terms does not change $\pi(\mathbf{x})$, but it gives rise to oscillatory modes that are filtered out here.
Importantly, a generalized DM method has been introduced that allows to obtain $\phi_{n}(x)$ also for non-gradient drift and position-dependent (multiplicative) noise [36], if one can estimate drift and diffusion at the data points. We implemented this generalized method for the high-dimensional ocean model Veros [50], showing that it gives very similar results compared to the simpler DM method (see App. A). While we do not consider examples with multiplicative noise, being able to properly account for it may be crucial in practice, as variance-based methods for EWS can fail [43, 44]. The advantage of using the original DM method is that it requires tuning fewer parameters, and it will be more widely applicable for EWS, since drift and diffusion matrix do not need to be estimated, which can be challenging for smaller data sets in high dimensions.
The DM algorithm defines a weighted graph on the data points, and subsequently computes the eigenvalues and eigenvectors of a random walk on this graph. To this end we define a kernel with bandwidth $\epsilon>0$ measuring the distance of two data points $\mathbf{x}$ and $\mathbf{y}$
@@
with $D_{i}=\sum_{j=1}^{N}\tilde{K}_{ij}$. The first few eigenvectors $\nu_{n}$ of $M$, corresponding to eigenvalues $\lambda_{n}^{(M)}$, define the so-called diffusion coordinates $\xi_{n}=\lambda_{n}^{(M)}\nu_{n}$. In the limit $N\to\infty$ and $\epsilon\to 0$, the operator $(M-I)/\epsilon$ converges to $\mathcal{L}^{*}$ (i.e. the backward FP operator) [31, 45, 46], with $\lambda_{n}^{(M)}=e^{\lambda_{n}}$. The Euclidean distance between data points in the DM coordinates is called the diffusion distance and measures how closely two points are connected via diffusion of the Markov chain $M$. Two points $\mathbf{x}$ and $\mathbf{y}$ may have small Euclidian distance, but large diffusion distance, which can reflect that the dynamics evolve on a lower dimensional manifold.
To evaluate the eigenfunctions approximately at points $\mathbf{y}$ not in the given data set $\{\mathbf{x}_{i}\}_{i=1}^{N}$, the so-called Nyström eigenspace interpolation can be employed
@@
and $D(\mathbf{y})=\sum_{i=1}^{N}\tilde{K}(\mathbf{y},\mathbf{x}_{i})$. In (14), $\xi_{n}(\mathbf{x}_{i})$ denotes the entry of the n-th eigenvector corresponding to the i-th data point.
In our DM implementation we normalize each data variable to have unit variance. Furthermore, we remove a small number of outliers so that we can use a smaller $\epsilon$ to obtain a better approximation of the backward eigenfunctions. Individual outliers are those data points that where the minimum of the distance to all other points is largest. We find $n=15$ points with the largest minimum distance and remove the respective columns and rows in the distance matrix $K(\mathbf{x}_{i},\mathbf{x}_{j})$ in (12). Thereafter we remove $n=10$ double outliers, which are those pairs of points where the second smallest distance to all other points is largest. The number of removed outliers has been chosen by trial and error to give the best performance across all data sets with typical sample size of about 10,000. If this step is skipped, often a quite large $\epsilon$ is needed to prevent the first diffusion coordinates from merely acting to cluster individual outliers against the rest of the data.